OpenStudy (anonymous):

The average speed of air molecules in your room is on the order of the speed of sound. What is their average velocity?

OpenStudy (aaronq):

\(\sf KE=\dfrac{3}{2}k_BT=\dfrac{1}{2}mv^2\) \(\sf v=\sqrt{\dfrac{3k_bT}{m}}=\sqrt{\dfrac{3k_bT*N_A}{M}}\) where \(\sf k_B\) is boltzman's constant, T is temp in Kelvin, M is molar mass of gas, \(\sf N_A\) is avogadro's number.